OpenStudy (kewlgeek555):

Can anyone be my tutor? I NEED someone good with explaining and solving recursive and explicit formulas. Please comment!

OpenStudy (kewlgeek555):

Please comment to let me know...

OpenStudy (kewlgeek555):

:-(

OpenStudy (dobby1):

I am prety good at math I could give it a try

OpenStudy (tkhunny):

It usually works better if you post a problem and include your best efforts at a solution.

OpenStudy (kewlgeek555):

Thank you @Dobby1 and @tkhunny I do not have a specific problem and I cannot think of one at the top of my head - I do not want answers I want explanations.

OpenStudy (dobby1):

oh uhm then i dont know because it helps if there is an actual problem I am helping with

OpenStudy (kewlgeek555):

@Dobby1 I am doing a quiz and I need to explain the steps on how I solved this recursive formula:\[f(1)=35, f(n)=f(1)+f(n-1)\ for\ n>1\]

OpenStudy (kewlgeek555):

That is a specific problem...

OpenStudy (dobby1):

ok let me take a look

OpenStudy (dobby1):

hm

OpenStudy (kewlgeek555):

That is not really the question - hold one, maybe this one is easier: Use the recursive formula that I gave you to make a table of values for 1 ≤ n ≤ 5. Show your calculations

OpenStudy (dobby1):

ok so that means if you had a number line n would be somewhere in between 1 and 5

OpenStudy (dobby1):

in other words n is no less than one but no greater than 5

OpenStudy (dobby1):

so a list of numbers that n could equal is 1 2 3 4 5

OpenStudy (dobby1):

does that make sense?

OpenStudy (kewlgeek555):

OMG, I am so sorry for not replying, I was afk.

OpenStudy (kewlgeek555):

So it would be 1, 2, 3, 4, 5?

OpenStudy (kewlgeek555):

@Dobby1

OpenStudy (anonymous):

These kind of questions are easy to do if you make a table: \[\begin{array}{c c c} n |& f(n)\\ 1 |& f(1) = 35 \\ 2 |& f(2) = f(1) + f(1) = 35 + 35 = 70\\ 3 |& f(3) = f(1) + f(2) = 35 + 70 = 105\\ 4 |& f(4) = f(1) + f(4-1) = ... \\ 5| & ... \end{array}\] and so on, taking previous values of f(n) from the table

OpenStudy (kewlgeek555):

So that would be the table, @Meepi ? Or is that an example?

OpenStudy (anonymous):

It's the same recursive formula as in the example, you just need to calculate it for n = 4 and n = 5 still :)

OpenStudy (kewlgeek555):

I meant is that a table for 1 is less than or equal to n is less than or equal to 5? I'm sorry, I didn't understand. But I will try to observe what you did so I can try to do it myself. :-D

OpenStudy (kewlgeek555):

kay, it is a table for that - sorry for getting confused, but let me see what 4 is.

OpenStudy (kewlgeek555):

So, based on what you did, @Meepi :\[35+105=140\]so...\[f(4)=f(1)+f(4-1)=140\]

OpenStudy (anonymous):

yep

OpenStudy (kewlgeek555):

Thanks. :-D